OpenStudy (anonymous):

f(x)=-x^2(x+7)(x^2-1) find the x and y intercept

OpenStudy (anonymous):

There are 5 x intercepts in this equation can you see any of them?

OpenStudy (anonymous):

is the y intercept -7

OpenStudy (anonymous):

You get the y intercept from setting x to zero, so y = -(0)^2*(0-7)*(0^2-1) So it would not be -7

OpenStudy (anonymous):

To get the x intercepts you have to set y = 0 in each different multiplication factor making up the polynomial, so: y = -x = 0 y = x = 0 y = x+7 = 0 y = x^2-1 = 0 You have to get x^2-1 into its factors The solution to that would look like (x+?) * (x-?) this way the middle term is canceled out

OpenStudy (mrnood):

for the x intercepts the function value =0 However it is already factorised for you so set each factor to 0 to get the solutions. (HINT - there are only 4 discrete solutions)

OpenStudy (jdoe0001):

http://www.mathsisfun.com/algebra/finding-intercepts-equation.html so, you set y=0 to find the x-intercepts and solve for "x" and x = 0 as shown above by Arfney and solve for "y"

OpenStudy (mrnood):

@Arfney I think you have a mistake (typo) in your post above: -x^2=0 x+7=0 x^2-1=0

OpenStudy (jdoe0001):

\(\large { f(x)=-x^2(x+7)(x^2-1) \\ \quad \\ \quad \\ y-intercept \\ \quad \\ f(x)=y=-{\color{brown}{ 0}}^2({\color{brown}{ 0}}+7)({\color{brown}{ 0}}^2-1) \\ \quad \\ x-intercept \\ \quad \\ {\color{brown}{ 0}}=-x^2(x+7)(x^2-1)\to \begin{cases} 0=x^2\\ 0=x+7\\ 0=x^2-1 \end{cases} }\)